OpenStudy (anonymous):

can you solve this? there is a parallelogram on the upper part : 2x-10 left part: 3y-4 bottom part : x+80 right part : y+20 ... whats the equation in x?

hartnn (hartnn):

hint : opposite sides of parallelogram are equal.

hartnn (hartnn):

upper part = bottom part

OpenStudy (anonymous):

so 2x-10=x+80 then the second equation is 3y-4=y+20? ryyt? :)

hartnn (hartnn):

absolutely! you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

hahahaha cool. i'm just making sure .... thankyou :))))

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

you're a genius :) hahaha .

hartnn (hartnn):

not really, but thanks for the compliment :D

OpenStudy (anonymous):

hahaha :))))) you really do like math huh? :)

hartnn (hartnn):

yes, very much. :)

hartnn (hartnn):

i see its your first day here, so \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

hahahaha thankyou so much & yea this is my 1st day :))